# On-Chain Analysis of Smart Contract Dependency Risks on Ethereum

### v1

@@ Abstract
In this paper, we present the first large-scale empirical study of smart contract dependencies, analyzing over 41 million contracts and 11 billion interactions on Ethereum up to December 2024. Our results yield four key insights: (1) 59% of contract transactions involve multiple contracts (median of 4 per transaction in 2024) indicating potential smart contract dependency risks; (2) the ecosystem exhibits extreme centralization, with just 11 (0.001%) deployers controlling 20.5 million (50%) of alive contracts, with major risks related to factory contracts and deployer privileges; (3) three most depended-upon contracts are mutable, meaning large parts of the ecosystem rely on contracts that can be altered at any time, which is a significant risk, (4) actual smart contract protocol dependencies are significantly more complex than officially documented, undermining Ethereum’s transparency ethos, and creating unnecessary attack surface. Our work provides the first large-scale empirical foundation for understanding smart contract dependency risks, offering crucial insights for developers, users, and security researchers in the blockchain space.
@@ I Introduction
First, we devise an original methodology to study smart contract on-chain dependencies through historical interactions. Our sound methodology is based on querying the state of the art of blockchain analytics platform. We design and implement queries to answer four research questions on the scale, centralization, popularity, and transparency of smart contracts.
We run our queries on the Ethereum blockchain, up to the last block of December 31, 2024. Overall, we analyze over 41 million contracts and 11 billion interactions between contracts. We analyze the results and their implications on security. Our large scale experiments yield the following findings.
@@
RQ3: Which smart contracts are most frequently called by other contracts? Over time, some smart contracts have become a cornerstone of the whole ecosystem. For those that are mutable, the major implication is that users’ funds are at risk if the smart contract owner is compromised.
@@
A sound methodology to study smart contract interactions and dependencies based on on-chain analysis. We use the state of the art blockchain analytics platform Allium.
@@ II Background
While numerous blockchain platforms exist [20], this paper focuses on Ethereum smart contracts. Ethereum is the most widely adopted smart contract platform powering a significant share of decentralized finance (DeFi), non-fungible tokens (NFTs), and other Web3 applications [49, 19]. Its wide adoption and mature ecosystem make Ethereum the ideal target for studying smart contract interactions at scale. This section introduces core concepts of Ethereum smart contracts, their deployment mechanisms, and their interaction via the Ethereum Virtual Machine (EVM).
@@ II-B Smart Contract Deployment
By Other Contracts (Factory Pattern): Contracts can programmatically deploy other contracts using the CREATE or CREATE2 opcodes. This allows for dynamic deployment of multiple contract instances with shared logic but varying initialization parameters.
@@ II-D Smart Contract Transparency
Unlike traditional software, smart contract source code must be made publicly available. The best practice is that smart contract developers push their smart contract source code on platforms like Etherscan [5] or Sourcify [12], a practice known as “verified smart contract”. Those platforms compile the provided source code to ensure that it matches the contract’s bytecode on the blockchain. Then, they show the verification status in UI and API, giving users confidence that the deployed contract functions as the ones for which source code is available [15, 41].
@@ III-A Research Questions
Transparency is a fundamental value proposition of blockchain technology, but the actual transparency of contract interactions may not live up to the expectations. Evaluating the gap between documented dependencies and on-chain reality helps users and stakeholders better assess risk when interacting with decentralized applications.
@@
We answer our research questions through on-chain analysis. Given the Ethereum mainnet’s enormous size (2 TB on a full node and 2.7 billion transactions), we must employ a specialized blockchain database that efficiently stores blockchain data in order to extract the information we need. This is called a blockchain analytics platform. To our knowledge, the three main ones are Allium, BigQuery, and Dune. We have compared them against ground truth data obtained from a self-hosted Erigon node. Our results indicate that Allium delivers the highest accuracy. For example, Allium is the only platform that includes interaction data of precompiled contracts which is required for our study. Besides, regarding the number of traces in 2024, the discrepancy between Allium and Erigon (ground truth) data is negligible, whereas BigQuery and Dune exhibited discrepancies higher than 7%. Consequently, in this paper, all numbers are extracted from Allium. With various queries, we examine the entire Ethereum mainnet dataset from the genesis block through the final block of 2024, spanning over 21 million blocks. All queries are made available on a replication repository at https://github.com/chains-project/crystal-clear.
@@ III-B Methodology
Distribution of contracts called per transaction: Let be the number of unique smart contracts involved within a single transaction . We analyze the distribution of for all transactions in a given time period. This metric provides insight into the degree of interconnectivity between contracts and the extent of their dependencies. To our knowledge, this aspect of smart contracts has never been studied before, previous work tends to only consider smart contracts in isolation (e.g. [30]).
@@
To determine the most widely interacted-with contracts, we analyze contract calls to identify the most called contracts on three non-depreacted call types as mentioned in section II-C.
These metrics allow us to assess how broadly a contract is integrated into the ecosystem, beyond just raw call counts. Contracts with a high number of calls are likely to serve as foundational component, widely relied upon across various dapps and applications.
This analysis also contributes to identify critical points of centralization (focal contracts). Centralization has potential reliability and security risks associated with a limited set of contracts.
@@ IV-A RQ1: Scale of Smart Contract Interactions
To understand the evolution of contract interactions, we measured the number of successful calls () each year and computed the proportions of each call type, namely , corresponding to regular calls, delegate calls, and static calls, respectively. Figure 3 illustrates both the annual evolution of and the distribution of call types. Over the past decade, the overall number of smart contract interactions has increased (from 1.56M to 2.59B), and also, the relative usage of call types has notably shifted. In the early years (2015–2017), regular calls dominated the interactions; however, starting in 2018, the usage of DELEGATECALL and STATICCALL began to increase, leading to a decline in the relative proportion of regular calls. By 2024, regular calls accounted for only 50% of all interactions. It is worth noting that in 2020, there was a pronounced surge in the usage of STATICCALL, it might be attributed to the popular DeFi protocol contracts starting to use it. Finally, the proportion of delegate calls also keeps increasing, up to 14.42% in 2024. This is explained by the popularity of the smart contract proxy pattern, and the need of dissociating data from behavior. Yet, it also poses security issues, we will elaborate on that point in RQ3.
[style=theoremstyle,frametitle=Answer to RQ1,linecolor=lightgray] Our results clearly demonstrate the presence of smart contract composition on Ethereum. In 2024, the median number of smart contracts involved in a transaction was 4. 4% of transactions involve more than 10 smart contracts and a few outliers involve up to thousand of smart contracts. To the best of our knowledge, our paper is the first to prove the reality of the vision of smart contract composition. This composition creates potential smart dependency risks, which is what we explore in the subsequent RQs.
@@ IV-B RQ2: Contract Creation
We analyze the dynamics of contract creation and its relationship to contract dependencies. With 41 million alive contracts on Ethereum, understanding who deploys them is essential. Figure 4 illustrates the entities which deploy them, with the top 100 contract deployers. Each deployer is represented as a bar, the height detailing the number of contracts deployed (left y-axis) and the red line being the cumulative contribution to the total alive contracts (right y-axis). First, we see that 83/100 top deployers are contracts themselves, with only 17 EOAs appearing in this top 100 list. This means that Ethereum heavily uses the contract factory pattern. The most prolific factory created 4.3M smart contracts (10% of all alive contracts). The top 11 contracts alone were responsible for deploying 50% of all alive contracts in 2024, and collectively, the top 100 deployers accounted for over 80% of all alive contracts. This shows that the large number of contracts does not reflect a large number of smart contract developers and smart contract operators. There is actually a handful of entities responsible for the majority of contract creation. This confirms related work on centralization in Ethereum [21].
@@
Table I provides a detailed overview of the top 10 deployers, including their human readable labels, source code availability (on Etherscan), their activity status (indicated by at least one transaction in December 2024, whether the created contracts’ behaviors are immutable (i.e., if a set of addresses has admin privilege to change the deployed contracts), the types of contracts they deploy, and the number of deployed contracts (same number as the corresponding bars on Figure 4. For instance, Cointool Batch minter [24] has deployed 4.3M contracts, its code is verified, actively used and mutable.
@@
It is positive to see that 6/10 contracts are verified, and surprising to see 4 unverified contracts responsible for creating 4.78M child contracts. Those contracts seem to be operated by crypto exchanges. Row “Active” clearly shows that some factories were heavily used in the past, yet have become dead since then, such as Bittrex (defunct company since December 2023) [1].
@@
Factory contracts, which deploy multiple instances from a single source, are attack vector, as a vulnerability in these factory contracts can spread to the child contracts. Table I shows that 8 out of these top 10 deployers primarily deploy proxy contracts, meaning that these child contracts interact with the implementation address using DELEGATECALL. 5 deployers implement EIP-1167 [3] proxies which means that all their created contracts are ”clones” and they redirect all their calls to one single address. This has key implications for security. If all child contracts are delegated to an address controlled by the deployer, this means that if the deployer is compromised, all child contracts become vulnerable as well, with their funds at risk. This is the case of Cointool Batch Minter and Bittrex Controller. For this reason, we strongly discourage this pattern. If the child contracts are immutable and do not give special privileged access to the deployer, the fact that a single factory contract creates many child contracts has no immediate implication for control-based security. This is the case for the 1inch CHI token and XENT token. However, even in such cases, immutability alone does not guarantee safety. If the cloned logic itself contains vulnerabilities, all child contracts would inherit those flaws, making a single attack successful on all child contracts at once. Thus, the widespread reuse of factory contracts amplifies the impact of vulnerabilities, regardless of upgradeability.
@@
OpenSea’s now deprecated Wyvern protocol is a prime example of how proxy and factory design patterns can introduce centralization risks. Under Wyvern, each user is assigned a personal proxy contract when they first interact with the marketplace. These proxies are programmed to delegate their function calls to a centralized logic contract, a single implementation contract controlled by a master key at Opensea. If this master key controlling the centralized implementation were to be compromised, an attacker could manipulate all user proxies at once, putting their NFTs and assets under existential or ownership risk.
@@
roxy risks are not theoretical. The recent Zoth exploit [47] confirms the real and immediate risks of centralized control in smart contract deployment. The deployer wallet of the protocol had admin privileges to the proxy contracts. Once compromised, a malicious delegate was deployed, leading to the loss of $8.4M.
@@ IV-C RQ3: Most Called Contracts
On Ethereum, the most called contracts are those which serve as fundamental building blocks for other contracts. We want to know what they are. We consider the four call types in EVM, but CALLCODE was deprecated in November 2018 and is now negligible, so we exclude it in the analysis of RQ3.
Figure 5 shows the result of most called contracts. First, we see that CALL, STATICCALL, and DELEGATECALL are all used in practice, they account for 28.71%, 41.31% and 31.03%, respectively. It means that the differences in their semantics have all found practical usage in successful DApp protocols.
We now look at CALL, the most standard call type, where the called contract maintains a separate state. Figure 5 shows the ten most called contracts with CALL. First, for this type of call, we see that 4 out of the 10 top callee contracts are tokens (WETH, USDC, USDT and Xen), confirming that tokens are fundamental building blocks. The most commonly called function is transfer.
The remaining 6 of 10 callee contracts are associated with decentralized finance. This empirically confirms that decentralized finance and programmable money a primary use case of smart contract composition. Because of the centrality of these contracts in the DeFi ecosystem, it is essential to pay special attention to their risk level. If one of them gets compromised, it may have contagion effects on the whole network of smart contracts.
@@
Out of the remaining two contracts called with STATICCALL: 1) the Uniswap Router contract used WETH() to return the address of the WETH, enabling accurate price retrieval. 2) the Chainlink Price Feed Contract is a blockchain oracle, a crucial component for obtaining external data for DApps. This distribution highlights the diversity of usages in STATICCALL interactions, confirming foundational infrastructure such as precompiled contracts and blockchain oracles for DeFi applications.
Finally, we look at DELEGATECALL, which represents 31.03% of the calls. DELEGATECALL enables to use the behavior of the callee contract on top of the state of the caller contract. It is essential for proxy contracts and upgrades, and also has other use cases. Five out of the 10 top callee contracts with DELEGATECALL are related to tokens. Four of them are implementation contracts of a proxy pattern (USDC, eETH, stETH, XEN Torrent), meaning that some tokens are implemented as proxies. Notably, 3 of them have been upgraded in the past. 3 times for USDC, 4 for eEth, 1 for stEth. For example, the USDC contract has been upgraded 3 times, progressing from version 1.0 to version 2.0, followed by versions 2.1 and 2.2. This is a major security concern: if the token maintainers get compromised, it would be possible to steal the entirety of the token funds. Overall, upgradable contracts are essential in the smart contract ecosystem of Ethereum, through DELEGATECALL interactions. The flexibility of upgrading contracts comes with the risk that any upgrade to these contracts can significantly influence the overall system integrity, if not destroy the upgradable contract and its dependents.
[style=theoremstyle,frametitle=Answer to RQ3,linecolor=lightgray] Token contracts dominate the standard calls on Ethereum, with 4 of the top 10 called contracts being tokens, empirically confirming programmable money as a primary use case. Notably, among contracts using DELEGATECALL, three out of every four token contracts have been upgraded in the past(e.g., USDC upgraded 3 times). This creates a systemic vulnerability: if token maintainers are compromised, entire token funds could be at risk. These findings demonstrate how certain contracts have become cornerstone infrastructure for the ecosystem, creating potential single points of failure with far-reaching implications for network security and stability.

### v2

@@ Abstract
In this paper, we present the first large-scale empirical study of smart contract dependencies by analyzing over 41 million contracts and 11 billion interactions on Ethereum up to December 2024. Our results yield four key insights: (1) 59% of contract transactions involve multiple contracts (median of 4 per transaction in 2024), indicating potential smart contract dependency risks; (2) the ecosystem exhibits extreme centralization, with just 11 (0.001%) deployers controlling 20.5 million (50%) of alive contracts, with major risks related to factory contracts and deployer privileges; (3) three most depended-upon contracts are mutable, meaning large parts of the ecosystem rely on contracts that can be altered at any time, which is a significant risk; (4) actual smart contract protocol dependencies are significantly more complex than officially documented, undermining Ethereum’s transparency ethos, and creating unnecessary attack surface. Our work provides the first large-scale empirical foundation for understanding smart contract dependency risks, offering crucial insights for developers, users, and security researchers in the blockchain space.
@@ I Introduction
First, we devise an original methodology to study smart contract on-chain dependencies through historical interactions. Our sound methodology is based on querying a state-of-the-art of blockchain analytics platform. We design and implement queries to answer four research questions on the scale, centralization, popularity, and transparency of smart contracts.
We run our queries on the Ethereum blockchain, up to the last block of December 31, 2024. Overall, we analyze over 41 million contracts and 11 billion interactions between contracts. We examine the results and their implications on security. Our large-scale experiments yield the following findings.
@@
RQ3: Which smart contracts are most frequently called by other contracts? Over time, some smart contracts have become a cornerstone of the whole ecosystem. For those that are mutable, the major implication is that users’ funds are at risk if the smart contracts’ owners are compromised.
@@
A sound methodology to study smart contract interactions and dependencies based on on-chain analysis. We use the state-of-the-art blockchain analytics platform Allium.
@@ II Background
While numerous blockchain platforms exist [20], this paper focuses on Ethereum smart contracts. Ethereum is the most widely adopted smart contract platform, powering a significant share of decentralized finance (DeFi), non-fungible tokens (NFTs), and other Web3 applications [48, 19]. Its wide adoption and mature ecosystem make Ethereum an ideal target for studying smart contract interactions at scale. This section introduces core concepts of Ethereum smart contracts, their deployment mechanisms, and their interaction via the Ethereum Virtual Machine (EVM).
@@ II-B Smart Contract Deployment
By Other Contracts (Factory Pattern): Contracts can programmatically deploy other contracts using the CREATE or CREATE2 opcodes. This allows for the dynamic deployment of multiple contract instances with shared logic but varying initialization parameters.
@@ II-D Smart Contract Transparency
Unlike traditional software, smart contract source code must be made publicly available. The best practice is that smart contract developers push their smart contract source code on platforms like Etherscan [5] or Sourcify [12], a practice known as “verified smart contract”. Those platforms compile the provided source code to ensure that it matches the contract’s bytecode on the blockchain. Then, they show the verification status in UI and API, giving users confidence that the deployed contract functions as the ones for which the source code is available [15, 40].
@@ III-A Research Questions
Transparency is a fundamental value proposition of blockchain technology, but the actual transparency of contract interactions may not live up to the expectations. Evaluating the gap between documented dependencies and on-chain reality helps users and stakeholders better to assess risks when interacting with decentralized applications.
@@
We answer our research questions through on-chain analysis. Given the Ethereum mainnet’s enormous size (2 TB on a full node and 2.7 billion transactions), we must employ a specialized blockchain database that efficiently stores blockchain data in order to extract the information we need. This is called a blockchain analytics platform. To our knowledge, the three main ones are Allium, BigQuery, and Dune. We have compared them against ground truth data obtained from a self-hosted Erigon node. Our results indicate that Allium delivers the highest accuracy. For example, Allium is the only platform that includes interaction data of precompiled contracts, which is required for our study. Besides, regarding the number of traces in 2024, the discrepancy between Allium and Erigon (ground truth) data is negligible, whereas BigQuery and Dune exhibited discrepancies higher than 7%. Consequently, in this paper, all numbers are extracted from Allium. With various queries, we examine the entire Ethereum mainnet dataset from the genesis block through the final block of 2024, spanning over 21 million blocks. All queries are made available on a replication repository at https://github.com/chains-project/crystal-clear.
@@ III-B Methodology
Distribution of contracts called per transaction: Let be the number of unique smart contracts involved within a single transaction . We analyze the distribution of for all transactions in a given time period. This metric provides insight into the degree of interconnectivity between contracts and the extent of their dependencies. To our knowledge, this aspect of smart contracts has never been studied before, as previous work tends to only consider smart contracts in isolation (e.g., [29]).
@@
To determine the most widely interacted-with contracts, we analyze contract calls to identify the most called contracts on three non-deprecated call types, as mentioned in section II-C.
These metrics allow us to assess how broadly a contract is integrated into the ecosystem, beyond just raw call counts. Contracts with a high number of calls are likely to serve as a foundational component, widely relied upon across various DApps and applications.
This analysis also contributes to identifying critical points of centralization (focal contracts). Centralization has potential reliability and security risks associated with a limited set of contracts.
@@ IV-A RQ1: Scale of Smart Contract Interactions
To understand the evolution of contract interactions, we measured the number of successful calls () each year and computed the proportions of each call type, namely , corresponding to regular calls, delegate calls, and static calls, respectively. Figure 3 illustrates both the annual evolution of and the distribution of call types. Over the past decade, the overall number of smart contract interactions has increased (from 1.56M to 2.59B), and also, the relative usage of call types has notably shifted. In the early years (2015–2017), regular calls dominated the interactions; however, starting in 2018, the usage of DELEGATECALL and STATICCALL began to increase, leading to a decline in the relative proportion of regular calls. By 2024, regular calls accounted for only 50% of all interactions. It is worth noting that in 2020, there was a pronounced surge in the usage of STATICCALL, which might be attributed to the popular DeFi protocol contracts starting to use it. Finally, the proportion of delegate calls also keeps increasing, up to 14.42% in 2024. This is explained by the popularity of the smart contract proxy pattern, and the need of dissociating data from behavior. Yet, it also poses security issues, which we will elaborate on that point in RQ3.
[style=theoremstyle,frametitle=Answer to RQ1,linecolor=lightgray] Our results clearly demonstrate the presence of smart contract composition on Ethereum. In 2024, the median number of smart contracts involved in a transaction was 4. 4% of transactions involve more than 10 smart contracts and a few outliers involve up to thousands of smart contracts. To the best of our knowledge, our paper is the first to prove the reality of the vision of smart contract composition. This composition creates potential smart dependency risks, which is what we explore in the subsequent RQs.
@@ IV-B RQ2: Contract Creation
We analyze the dynamics of contract creation and its relationship to contract dependencies. With 41 million alive contracts on Ethereum, understanding who deploys them is essential. Figure 4 illustrates the entities which deploy them, with the top 100 contract deployers. Each deployer is represented as a bar, the height detailing the number of contracts deployed (left y-axis) and the red line being the cumulative contribution to the total alive contracts (right y-axis). First, we see that 83/100 top deployers are contracts themselves, with only 17 EOAs appearing in the top 100 list. This means that Ethereum heavily uses the contract factory pattern. The most prolific factory created 4.3M smart contracts (10% of all alive contracts). The top 11 contracts alone were responsible for deploying 50% of all alive contracts in 2024, and collectively, the top 100 deployers accounted for over 80% of all alive contracts. This shows that the large number of contracts does not reflect a large number of smart contract developers and smart contract operators. There is actually a handful of entities responsible for the majority of contract creation. This confirms related work on centralization in Ethereum [21].
@@
Table I provides a detailed overview of the top 10 deployers, including their human readable labels, source code availability (on Etherscan), activity status (indicated by at least one transaction in December 2024, whether the created contracts’ behaviors are immutable (i.e., if a set of addresses has admin privilege to change the deployed contracts), the types of contracts they deploy, and the number of deployed contracts (same number as the corresponding bars on Figure 4. For instance, Cointool BatchMinter [24] has deployed 4.3M contracts, its code is verified, actively used and mutable.
@@
It is positive to see that 6/10 contracts are verified, and surprising to see 4 unverified contracts responsible for creating 4.78M child contracts. Those contracts seem to be operated by crypto exchanges. The Active row clearly shows that some factories were heavily used in the past, yet have become dead since then, such as Bittrex (defunct company since December 2023) [1].
@@
Factory contracts, which deploy multiple instances from a single source, are attack vectors, as a vulnerability in them can spread to child contracts. Table I shows that 8 out of these top 10 deployers primarily deploy proxy contracts, meaning that these child contracts interact with the implementation address using DELEGATECALL. 5 deployers implement EIP-1167 [3] proxies which means that all their created contracts are “clones” and they redirect all their calls to one single address. This has key implications for security. If all child contracts are delegated to an address controlled by the deployer, this means that if the deployer is compromised, all child contracts become vulnerable as well, with their funds at risk. This is the case of Cointool BatchMinter and Bittrex Controller. For this reason, we strongly discourage this pattern. If the child contracts are immutable and do not give special privileged access to the deployer, the fact that a single factory contract creates many child contracts has no immediate implication for control-based security. This is the case for the 1inch CHI token and XENT token. However, even in such cases, immutability alone does not guarantee safety. If the cloned logic itself contains vulnerabilities, all child contracts would inherit those flaws, making a single attack successful on all child contracts at once. Thus, the widespread reuse of factory contracts amplifies the impact of vulnerabilities, regardless of upgradeability.
@@
OpenSea’s now deprecated Wyvern protocol is a prime example of how proxy and factory design patterns can introduce centralization risks. Under Wyvern, each user is assigned a personal proxy contract when they first interact with the marketplace. These proxies are programmed to delegate their function calls to a centralized logic contract, a single implementation contract controlled by a master key at Opensea. If this master key which controls the centralized implementation, were to be compromised, an attacker could manipulate all user proxies at once, putting their NFTs and assets under existential or ownership risk.
@@
Proxy risks are not theoretical. The recent Zoth exploit [46] confirms the real and immediate risks of centralized control in smart contract deployment. The deployer wallet of the protocol had admin privileges to the proxy contracts. Once compromised, a malicious delegate was deployed, leading to the loss of $8.4M.
@@ IV-C RQ3: Most Called Contracts
On Ethereum, the most called contracts are those that serve as fundamental building blocks for other contracts. We want to know what they are. We consider the four call types in EVM, but CALLCODE was deprecated in November 2018 and is now negligible, so we exclude it in the analysis of RQ3.
Figure 5 shows the result of most frequently called contracts. First, we see that CALL, STATICCALL, and DELEGATECALL are all used in practice; they account for 28.71%, 41.31% and 31.03% of the calls, respectively. It means that the differences in their semantics have all found practical usage in successful DApp protocols.
We now look at CALL, the most standard call type, where the called contract maintains a separate state. Figure 5 shows the ten most called contracts with CALL. First, for this type of call, we see that 4 out of the 10 top callee contracts are tokens (WETH, USDC, USDT, and Xen), confirming that tokens are fundamental building blocks of Ethereum, with transfer being the most frequently called function.
The remaining 6 out of 10 callee contracts are associated with decentralized finance. This empirically confirms that decentralized finance and programmable money a primary use cases of smart contract composition. Because of the centrality of these contracts in the DeFi ecosystem, it is essential to pay special attention to their risk level. If one of them gets compromised, it may have contagion effects on the whole network of smart contracts.
@@
Out of the remaining two contracts called with STATICCALL: 1) the Uniswap Router contract used WETH() to return the address of the WETH, enabling accurate price retrieval. 2) the Chainlink Price Feed Contract is a blockchain oracle, a crucial component for obtaining external data for DApps. This distribution highlights the diversity of usages in STATICCALL interactions, confirming foundational infrastructures such as precompiled contracts and blockchain oracles for DeFi applications.
Finally, we look at DELEGATECALL, which represents 31.03% of the calls. DELEGATECALL enables to use the behavior of the callee contract on top of the state of the caller contract. It is essential for proxy contracts and upgrades, and also has other use cases. 5 out of the 10 top callee contracts with DELEGATECALL are related to tokens. 4 of them are implementation contracts of a proxy pattern (USDC, eETH, stETH, XEN Torrent), meaning that some tokens are implemented as proxies. Notably, 3 of them have been upgraded in the past. 3 times for USDC, 4 for eEth, and 1 for stEth. For example, the USDC contract has been upgraded 3 times, progressing from version 1.0 to version 2.0, followed by versions 2.1 and 2.2. This is a major security concern: if the token maintainers get compromised, it would be possible to steal the entirety of the token funds. Overall, upgradable contracts are essential in the smart contract ecosystem of Ethereum, through DELEGATECALL interactions. The flexibility of upgrading contracts comes with the risk that any upgrade to these contracts can significantly influence the overall system integrity, if not destroy the upgradable contract and its dependents.
[style=theoremstyle,frametitle=Answer to RQ3,linecolor=lightgray] Token contracts dominate the standard calls on Ethereum, with 4 of the top 10 called contracts being tokens, empirically confirming programmable money as the primary use case. Notably, among contracts using DELEGATECALL, 3 out of 4 token contracts have been upgraded in the past (e.g., USDC upgraded 3 times). This creates a systemic vulnerability: if token maintainers are compromised, entire token funds could be at risk. These findings demonstrate how certain contracts have become cornerstone infrastructure for the ecosystem, creating potential single points of failure with far-reaching implications for network security and stability.